OpenStudy (anonymous):

Graph the solution to the following inequality on the number line. (x^2)-2x<=8

OpenStudy (anonymous):

\[x ^{2}-2x \le8\]

OpenStudy (shamil98):

Do you know the quadratic formula? You can use that to find the solutions

OpenStudy (anonymous):

quadratic ok working..

OpenStudy (shamil98):

x^2 -2x - 8 <= 0

OpenStudy (campbell_st):

ok... so you have x^2 - 2x - 8 <= 0 can you factor it...?

OpenStudy (shamil98):

yeah you can factor this as well

OpenStudy (anonymous):

yea i think i should factor it

OpenStudy (anonymous):

and then test

OpenStudy (anonymous):

i remember thanks team

OpenStudy (campbell_st):

or you can use complete the square \[x^2 -2x + c \le 8\] may be an easier way to get a solution...

OpenStudy (shamil98):

Just to clarify, this is what you would use to graph it from. (x-4)(x+2) <= 0 -2 <= x <= 4

OpenStudy (shamil98):

Since it has the or equal to sign it's going to be a closed circle

OpenStudy (campbell_st):

there you go... no need to test when you've been given the solution

OpenStudy (anonymous):

sweet thats what i got i forgot i had to set to zero and find my zeros

OpenStudy (shamil98):

Then just plot it on a number line

OpenStudy (anonymous):

roger that